# The New LLM Bottleneck: A Systems Perspective on Latent Attention and Mixture-of-Experts

### v1

@@ II-A Standard LLM architecture and its layers
For the FC layers, all requests share the same weights in LLM inference. Batching multiple requests enables the FC weight reuse, reducing memory access overhead. As memory bandwidth scales more slowly than compute capability in modern accelerators (Table I), it becomes a major bottleneck, especially in the decode stage, where GEMV operations offer limited data reuse. While the batching improves throughput, it does not reduce per-request latency, and the batch size $B$ is bounded by SLOs [59]. Moreover, as each request maintains its own KV\$, a larger $B$ increases memory capacity requirements, making memory capacity another key limiter for the maximum feasible value.
@@ II-D LLM serving systems
Pipeline parallelism (PP), an inter-layer model parallelization technique, partitions sequential layers into multiple accelerators to address memory capacity limitations with lower communication cost over TP. However, TP is often preferred over PP in LLM inference as it directly reduces the latency of computations, whereas PP can not. Another core weakness of PP is the creation of inevitable bubbles, where accelerators sit idle as they have to wait for the input data from the previous layer computed by other accelerators in the pipeline. Although it can be mitigated by processing several micro-batches to keep the pipeline full, this method’s effectiveness diminishes with smaller batch sizes, leading to hardware under-utilization. Thus, in this paper, we focus on TP rather than PP.
@@ IV Demystifying the Compound Interplay of Algorithm–Hardware–Parallelism
Figure 3: Time per output token and per-device throughput of DeepSeek-R1 and GPT-3 across varying sequence length $L$ and batch size $B$. The blurred area indicates configurations where out-of-memory errors occurred and the red dashed line represents the time required to read all data from HBM, which determines the maximum feasible combinations of $(B,L)$. The experiment assumes a 32 NVIDIA B200 GPU system.
@@ IV-A The latest LLM wears MLA and MoE
As the sequence length and batch size ($B$) increase, core attention becomes the main bottleneck in the decode stage, limiting throughput and leaving hardware underutilized [39, 56, 20]. The maximum feasible $B$ in an LLM system is determined by the memory capacity that remains after storing the model parameters because the residual capacity accommodates the per-request KV\$. FC layers that use weights are in a memory-bound region when $B$ is smaller than $RP_{\mathrm{acc}}$. With an insufficient maximum $B$, time per output token (TPOT) is bounded by the time required to read the entire memory, leaving computing resources in an accelerator idle.
@@ V Insights on Attention
with $B2n_{\mathrm{hd}}Ld_{\mathrm{KVco}}$ $2B(d_{\mathrm{KVco}}L+n_{\mathrm{hd}}L)$ $\approx(d_{\mathrm{hd}}^{-1}+d_{\mathrm{KVco}}^{-1})^{-1}$ $\approx 100$
@@
Figure 5: (a) Normalized latency of the attention block in the decode stage without reordering compared to a reordered Attention block. (b) and (c) shows execution time ratio of each layer in the attention block in the decode stage with and without reordering, across varying sequence length $L$ and batch size $B$. The experiments assume a 32 NVIDIA B200 GPU system.
@@ V-B Impact of Reordering MLA
Without reordering, the core attention in MLA preserves the same computational flow as MHA except for the runtime compression and decompression steps immediately preceding it. The required computation (FLOPs) and memory access for the K decompression and score layers are proportional to $B$ and $L$ during the decode stage (see Table III), whereas those of the FC layers do not scale with $L$. Thus, KV decompression and core attention dominate the execution time of an attention block as $L$ increases (see Figure 5(b)). At $B\!=\!256$ and $L\!=\!4096$, KV decompression and core attention account for 59% and 40% of an attention block latency, respectively.
@@
By contrast, this benefit disappears in the prefill stage as the size of $\mathbf{Q}$ is proportional to $L\!=\!L_{\mathrm{in}}$, making the computational cost of K decompression unchanged. The reordering rather increases the amount of computation required for the score layer by $\nicefrac{{d_{\mathrm{KVco}}}}{{d_{\mathrm{hd}}}}$ (4 for DeepSeek-R1) times in both stages, as it replaces the original score layer with a multiplication between $\mathbf{Q}_{i}\cdot{\mathbf{W}_{\mathrm{DK_{i}}}}^{\mathrm{T}}$ and ${\mathbf{C}_{\mathrm{KV}}}^{\mathrm{T}}$ (see Eq. 7), where one of the matrix dimensions becomes $d_{\mathrm{KVco}}$ instead of $d_{\mathrm{hd}}$.
@@
FlashMLA [23], a GPU-optimized implementation, further doubles this Op/B by reusing ${\mathbf{C}_{\mathrm{KV}}}$ loaded during the score layer in the subsequent context layer. A $2n_{\mathrm{hd}}$ Op/B (i.e., 256 in DeepSeek-R1) closely approaches the ridge point of modern accelerators (Obs.S2). Since the a.i. of core attention is near $RP_{\mathrm{acc}}$ of modern accelerators, its performance is balanced between compute-bound and memory-bound. This means the time spent on computation is still approximately equivalent to the time spent on $\mathbf{C}_{\mathrm{KV}}$ access, thereby the throughput of core attention improved approximately by $2d_{\mathrm{dec}}/d_{\mathrm{KVco}}$ ($=\!64$ for DeepSeek-R1).
@@
Putting it all together, layer reordering significantly reduces the latency of KV decompression and core attention within the overall attention block latency in the decode stage (see Figure 5(a)). Thus, layer reordering improves the execution time of the attention block in the decode stage by 119.07$\times$, when $B\!=\!256$ and $L\!=\!4096$. When $B\!=\!1$ and $L\!=\!4096$, the attention block’s latency in the prefill stage with reordering is 2.02$\times$ higher than that without reordering. Hereafter, the prefill stage uses MLA without reordering and the decode stage uses MLA with reordering.
Observation-P3: Layer reordering enhances the attention block’s throughput by substantially improving its a.i., primarily by reducing memory access despite the additional computational overhead in the decode stage; however, it degrades performance in the prefill stage.
MLA diminishes the need for specialized attention hardware. In contrast to MHA or GQA, where core attention in decode stages typically shows low a.i. (Op/B in the 1–8 range) [56], prompting the development of processing-in-memory or near-data-processing architectures [39, 56, 20, 28] for acceleration, MLA and reordering significantly increase the Op/B of core attention, approaching $\mathrm{RP_{acc}}$. Consequently, previous attention-specific hardware is less effective for accelerating MLA attention because performance is no longer constrained by memory bandwidth.
@@ V-C Parallelism on MLA
Figure 7: Latency of the Attention block with and without reordering as batch sizes and $\mathrm{deg}_{\mathrm{TP}}$ vary when $L\!=\!4096$.
@@ VII-A The Synergistic impact of MLA and MoE
Figure 10: System throughput and execution time ratio of operations in decode stage of DeepSeek-R1, varying batch sizes and sequence lengths. In this experiment, we assume a 64 B200 GPU system.
@@ VII-B Scaling the system out
Figure 11: System throughput and execution time ratio of decode stage of DeepSeek-R1 when using InfiniBand XDR (800Gb/s) among a group of GPUs (DGX) varying batch sizes and sequence length. We assume 32 B200 GPU system.

### v2

@@ II-A Standard LLM architecture and its layers
For the FC layers, all requests share the same weights in LLM inference. Batching multiple requests enables the FC weight reuse, reducing memory access overhead. As memory bandwidth scales more slowly than compute capability in modern accelerators (Table I), it becomes a major bottleneck, especially in the decode stage, where GEMV operations offer limited data reuse. While the batching improves throughput, it does not reduce per-request latency, and the system-wise batch size $B$ is bounded by SLOs [58]. Moreover, as each request maintains its own KV\$, a larger $B$ increases memory capacity requirements, making memory capacity another key limiter for the maximum feasible value.
@@ II-D LLM serving systems
We do not adopt pipeline parallelism (PP) for LLM inference due to its limited ability to reduce computation latency and poor hardware utilization. While PP mitigates memory constraints with lower communication overhead than TP, it suffers from pipeline bubbles. Though micro-batching can reduce these bubbles, its effectiveness diminishes with smaller batches.
@@ IV Demystifying the Compound Interplay of Algorithm–Hardware–Parallelism
Figure 3: Time per output token and per-device throughput of DeepSeek-R1 and GPT-3 across varying sequence length and batch size. The blurred area indicates configurations where out-of-memory errors occurred and the red dashed line represents the time required to read all data from HBM, which determines the maximum feasible combinations of sequence length and batch size. The experiment assumes a 32 NVIDIA B200 GPU system.
@@ IV-A The latest LLM wears MLA and MoE
As the sequence length ($L$) and batch size ($B$) increase, core attention becomes the main bottleneck in the decode stage, limiting throughput and leaving hardware underutilized [38, 55, 20]. The maximum feasible $B$ in an LLM system is determined by the memory capacity that remains after storing the model parameters because the residual capacity accommodates the per-request KV\$. FC layers that use weights are in a memory-bound region when $B$ is smaller than $RP_{\mathrm{acc}}$. With an insufficient maximum $B$, time per output token (TPOT) is bounded by the time required to read the entire memory, leaving computing resources in an accelerator idle.
@@ V Insights on Attention
with $B2n_{\mathrm{hd}}Ld_{\mathrm{KVco}}$ $2B(d_{\mathrm{KVco}}L+n_{\mathrm{hd}}L)$ $\approx(n_{\mathrm{hd}}^{-1}+d_{\mathrm{KVco}}^{-1})^{-1}$ $\approx 100$
@@
Figure 5: (a) Normalized latency of the attention block in the decode stage without reordering compared to a reordered Attention block. (b) and (c) shows execution time ratio of each layer in the attention block in the decode stage with and without reordering, across varying sequence length and batch size. The experiments assume a 32 NVIDIA B200 GPU system.
@@ V-B Impact of Reordering MLA
Without reordering, the core attention in MLA preserves the same computational flow as MHA except for the runtime compression and decompression steps immediately preceding it. The required computation (FLOPs) and memory access for the K decompression and score layers are proportional to $B$ and $L$ during the decode stage (see Table III), whereas those of the FC layers do not scale with $L$. Thus, KV decompression and core attention dominate the execution time of an attention block as $L$ increases (see Figure 5(b)). At $B\!=\!128$ and $L\!=\!4096$, KV decompression and core attention account for 59% and 40% of an attention block latency, respectively.
@@
By contrast, this benefit disappears in the prefill stage as the size of $\mathbf{Q}$ is proportional to $L\!=\!L_{\mathrm{in}}$, making the computational cost of K decompression unchanged. The reordering rather increases the amount of computation required for the score layer by $\nicefrac{{d_{\mathrm{KVco}}}}{{d_{\mathrm{hd}}}}$ (4 for DeepSeek-R1) times in both stages, as it replaces the original score layer with a multiplication between $\mathbf{Q}_{i}\cdot{\mathbf{W}_{\mathrm{DK_{i}}}}^{\mathrm{T}}$ and ${\mathbf{C}_{\mathrm{KV}}}^{\mathrm{T}}$ (see Eq. 7), where dimension of each head becomes $d_{\mathrm{KVco}}$ instead of $d_{\mathrm{hd}}$.
@@
FlashMLA [23], a GPU-optimized implementation, further doubles this Op/B by reusing ${\mathbf{C}_{\mathrm{KV}}}$ loaded during the score layer in the subsequent context layer. A $2n_{\mathrm{hd}}d_{\mathrm{KVco}}/(n_{\mathrm{hd}}+d_{\mathrm{KVco}})$ Op/B (i.e., $\approx$ 200 in DeepSeek-R1) closely approaches the ridge point of modern accelerators (Obs.S2). Since the a.i. of core attention is near $RP_{\mathrm{acc}}$ of modern accelerators, its performance is balanced between compute-bound and memory-bound. Consequently the time spent on computation remains approximately equal to the time spent on $\mathbf{C}_{\mathrm{KV}}$ access, thereby the latency of core attention reduced approximately by $2d_{\mathrm{dec}}/d_{\mathrm{KVco}}$ ($=\!64$ for DeepSeek-R1) after layer reordering.
Despite the use of latent space, explicitly generating decompressed $\mathbf{K}$ and $\mathbf{V}$ requires a significant amount of memory for activation, limiting $B$ as shown in Figure 5. For DeepSeek-R1, decompressing the $\mathbf{K}$ tensor at a per-accelerator batch size of 256 and $L=4096$, inflates the activation footprint to $\approx$50GB. As layer reordering shrinks the size of on-the-fly activation, the maximum feasible $B$ increased and delivers proportionally higher throughput on FC layers.
@@
Putting it all together, layer reordering significantly reduces the latency of the attention block by up to 103.12$\times$ (see Figure 5. When batch size per accelerator $=\!1$ and $L\!=\!4096$, the attention block’s latency in the prefill stage with reordering is 2.02$\times$ higher than that without reordering. Hereafter, the prefill stage uses MLA without reordering and the decode stage uses MLA with reordering.
Observation-P3: Layer reordering substantially reduces the latency of attention block by compressing the size of KV cache to be read and drastically mitigating the impact of KV decompression overhead.
@@ V-C Parallelism on MLA
Figure 7: Latency of the Attention block with and without reordering as batch sizes and $\mathrm{deg}_{\mathrm{TP}}=n_{\mathrm{acc}}$ vary when $L\!=\!4096$.
@@ VII-A The Synergistic impact of MLA and MoE
Figure 10: System throughput and execution time ratio of operations in decode stage of DeepSeek-R1, varying sequence lengths and batch sizes. In this experiment, we assume a 64 B200 GPU system.
@@ VII-B Scaling the system out
Figure 11: System throughput and execution time ratio of decode stage of DeepSeek-R1 when using InfiniBand XDR (800Gb/s) among a group of GPUs (DGX) varying sequence lengths and batch sizes. We assume 32 B200 GPU system.